# FedSynthCT-Brain: A Federated Learning Framework for Multi-Institutional Brain MRI-to-CT Synthesis

### v1

@@ Abstract
The generation of Synthetic Computed Tomography (sCT) images has become a pivotal methodology in modern clinical practice, particularly in the context of Radiotherapy (RT) treatment planning. The use of sCT enables the calculation of doses, pushing towards Magnetic Resonance Imaging (MRI) guided radiotherapy treatments. Moreover, with the introduction of MRI-Positron Emission Tomography (PET) hybrid scanners, the derivation of sCT from MRI can improve the attenuation correction of PET images. Deep learning methods for MRI-to-sCT have shown promising results, but their reliance on single-centre training dataset limits generalisation capabilities to diverse clinical settings. Moreover, creating centralised multicentre datasets may pose privacy concerns. To solve the issues, this study introduces FedSynthCT-Brain, a framework based on the Federated Learning (FL) paradigm for MRI-to-sCT in brain imaging. We reproduced a federation across four European and American centres using a U-Net-based model. The approach was implemented using data from centres belonging the federation and it was tested on an unseen dataset from a centre outside the federation. In the case of the unseen centre, the federated model achieved a median Mean Absolute Error (MAE) of $102.0$ HU across 23 patients, with an interquartile range of $96.7-110.5$ HU. The median (interquartile range) for the Structural Similarity Index (SSIM) and the Peak Signal to Noise Ratio (PNSR) were $0.89$ $(0.86-0.89)$ and $26.58$ $(25.52-27.42)$, respectively. The analysis of the results showed acceptable performances of the federated approach, thus highlighting the potential of FL to enhance MRI-to-sCT to improve generalisability and advancing safe and equitable clinical applications while fostering collaboration and preserving data privacy.
@@ 1 Introduction
Intermodality medical Image-To-Image (I2I) translation refers to the process of mapping images from one modality (e.g., Magnetic Resonance Imaging - MRI, Computed Tomography - CT, Positron Emission Tomography - PET) to another, while preserving essential structural and anatomical information. Deep Learning (DL) based techniques have been shown to outperform traditional methods based on image registration and/or voxel-wise analysis [10] and they have gained attention in several clinical fields, the most important being the derivation of Synthetic CT (sCT) from MRI or Cone Beam CT (CBCT), to facilitate radiotherapy planning/adaptation or to improve attenuation correction in PET/MRI acquisition [49, 52].
@@
To improve generalizability in MRI-to-sCT and foster innovation in the field, initiatives like the SynthRAD2023 [21] challenge have been proposed by the scientific community. These efforts allow DL models — such as the MSEP framework, the challenge winner [20, 8] — to be trained on multi-centre datasets, providing access to more diverse patient data. While directly pooling medical images from multiple sites could intuitively address the issue of small sample sizes, this approach faces significant barriers due to stringent privacy protection policies such as the Health Insurance Portability and Accountability Act (HIPAA)  [^1] https://www.hhs.gov/hipaa/index.html in the United States and the General Data Protection Regulation (GDPR) [^2] https://gdpr-info.eu/ in Europe. These regulations impose limitations on the exchange of personal health data, posing practical challenging to sharing and centralising medical images across sites [23, 16].
@@
To the best of the authors’ knowledge, this is among the first application in literature to employ a FL paradigm in MRI-to-sCT. The main contributions of the work are:
@@
  • We validate the proposed approach on real multi-centre data, simulating an authentic scenario where each centre handles heterogeneous data, due to different scanners and acquisition modalities.
  • We evaluate the proposed FedSynthCT-Brain on an independent dataset outside the federation to assess its generalization capability on data from external centres, demonstrating its effectiveness even in the presence of significant data heterogeneity.
@@ 2.1 Centralised MRI-to-sCT
Over the past years, numerous DL methods have been implemented for MRI-to-sCT. The proposed model architectures and methodologies have become increasingly complex in order to achieve accurate results. Among the most widely used models are: U-Nets, Generative Adversarial Networks (GANs), Transformers, and Diffusion-based architectures [10].
Several works employed three different 2-dimensional (2D) fully Convolutional Neural Networks (CNN), one for each anatomical plane (axial, coronal, and sagittal), to retain structural information in sCTs while simultaneously reducing computational costs, especially when using U-Nets [10, 28, 50, 19]. These 2D methods, by processing slices independently or in parallel, inherently simplify training in distributed environments, such as FL, where data sharing is constrained by computational heterogeneity [3].
On the other side, 3D-DL approaches have demonstrated improved spatial continuity and consistency in terms of Hounsfield Unit (HU) across slices compared to their 2D akin [10]. However, the primary limitation of 3D models lies in their higher computational demand, which requires splitting volumetric data into smaller 3D patches for processing [26, 25]. This patch-based processing may lead to the loss of global context and structural coherence, especially when patches are analysed in isolation. Moreover, the scalability of 3D approaches in FL settings is further constrained due to the increased communication overhead and memory requirements associated with volumetric data. These factors make 3D-DL approaches less suitable for FL real-world scenarios where limited computational resources requirements must be accounted for [3].
@@ 2.2 Federated learning in medical imaging
More recently, FL approaches were proposed for segmentation tasks, such as for brain tumour [53, 2], prostate [61], optic cup and disc segmentation [43, 55]. This led to the development of frameworks that integrate the aforementioned works and are capable of performing segmentations of multiple anatomical structures or different image modalities using FL and foundational models, such as the FedFMS [38].
Recent advancements in medical image synthesis have introduced innovative methods, such as FedMed-GAN, which specifically addresses the challenge of synthesizing cross-modality brain MRI images in an FL setting [56]. It employs FedAvg as the aggregation strategy, enabling the combination of locally trained models from multiple institutions.
Inspired by the FL paradigm, this study explores the potential of FL for MRI-to-CT synthesis, tackling a challenge that remains unaddressed in existing research.
@@ 3.1 Datasets and preprocessing
Data origin Patient ##\## Image modality - Voxel size Size
[mm3] Axial Sagittal Coronal
Centre A (USA) 15 MR - 1 $\times$ 1 $\times$ 1 CT - 0.49 - 0.67 $\times$ 0.49 - 0.67 $\times$ 2.5 256 176 256
Centre B (IT) 14 MR - 1 $\times$ 1 $\times$ 1 CT - 0.98 $\times$ 0.98 $\times$ 3.27 256 176 248
Centre C (DE) 21 MR - 0.78 $\times$ 0.78 $\times$ 1 CT - 0.78 $\times$ 0.78 $\times$ 1 226 - 357 512 512
Centre D (NL, SynthRAD Centre C [21]) 29 MR - 0.98 – 1.12 $\times$ 0.98 – 1.12 $\times$ 0.98 – 1.12 CT - 0.69 – 0.78 $\times$ 0.69 – 0.79 $\times$ 1 - 3 167 - 213 216 - 262 250 - 277
Centre E (NL, SynthRAD Centre B [21]) 23 MR - 0.98 $\times$ 0.98 $\times$ 0.98 CT - 0.59 - 1.27 $\times$ 0.59 - 1.27 $\times$ 1 - 2 167 - 262 200 - 225 225 - 248
Data origin MRI scanners / Field strength [T] CT scanners / kVp
Centre A (USA) MAGNETOM Trio / 3 LightSpeed QX/i / 140
Centre B (IT) Biograph mMR / 3 Discovery ST / 120
Centre C (DE) Vantage Titan / 1.5 Brilliance Sensation Open / 120
Centre D (NL, SynthRAD Centre C [21]) MAGNETOM, Avanto_fit, Skyra, Vida_fit, Prisma_fit / 1.5-3 Brilliance Big Bore / 120
Centre E (NL, SynthRAD Centre B [21]) MAGNETOM Aera, Avanto_fit / 1.5-3 SOMATOM Definition AS / 120
@@
MRIs from Centre B were acquired using a PET-MRI hybrid scanner. Differently from the other centres, this acquisitions were not performed for radiotherapy planning purposes. Thus, the tube energy in the CT was 120kV and the slice thickness was 3.75 mm which introduced partial volume artifacts in the upper part of the skull bone. In light of the substantial difference in the acquisition protocol and scanners, Centre B dataset was used to increase the image variability within the federation (i.e. to introduce noise) to test the robustness of the federated model and simulate a real context.
Images belonging to Centre C presented higher spatial resolution in the cranial caudal direction compared to the other datasets. However, the Centre C’s dataset presented a number of additional challenges, including: the absence of the top of the skull in some patients, the presence of the scanner bed and pillow for each CT, metal artefacts, and high bias field artifact.
The Centre D’s and Centre E’s datasets derived exclusively from two different institutions of the SynthRAD2023 multi-centre dataset (SynthRAD Centre C and SynthRAD Centre B respectively), which was created for the SynthRAD challenge within different MRI-to-sCT algorithms [21]. Before publication, the datasets were cropped to remove the facial region. TheFOV covered by both imaging modalities was similar, except for a cut-off at the lower end with a slight angle in addition to a straight horizontal line. In general, Center D and E presented less bias field artifact when compared to Center C. However, between Centre D and E, it was more pronounced in the Centre E dataset. Furthermore, both Centre MRIs were obtained using gadolinium-based contrast agents, thus introducing further heterogeneity within the federation.
To address the image variability, a pre-processing strategy was implemented at each client site with the aim of harmonising data within the federation, without the exchange of data. Following [54], N4 Bias Field Correction was applied to the MRI data to correct intensity inhomogeneity, which can result from factors such as magnetic field variations or patient positioning. Additionally, Min-Max Normalization was applied within each client to reduce variability in pixel intensity values.
@@ 3.3 Deep learning model
To train the aforementioned architecture, each client input 2D slices from each MRI in 3 anatomical planes, axial (Ax), coronal (Cor), and sagittal (Sag), as it is shown in Figure 4. Slices were presented to the network in a randomised order. After extracting all slices from each anatomical plane, they were shuffled to ensure that the DL model received non-sequential input from more than one anatomical plane at a time. This minimised the risk of the model becoming overly dependent on any single plane or orientation, encouraging comprehensive and plane-invariant feature learning across all planes [24].
@@ 4.3 Performance metrics
In order to evaluate the efficacy of the federated model both on the clients test datasets (centres A, B, C, and D), and unseen test dataset (centre E) standard image similarity metrics in the MRI to CT translation task were employed [49]. These included:
@@ 4.4.1 Deep learning architecture selection
We explored several variants of the U-Net architecture which has been used for centralised MRI-to-sCT-Simple U-Net [45], Spadea et al. U-Net [51], Fu et al. U-Net [13], Li et al. U-Nets [30]. These variants included variations in implementations, such as changes in network depth, feature map sizes, and the integration of attention mechanisms. Our analysis focused on evaluating the balance between computational complexity and generalization capability across these architectures.
Despite its inherent complexity, we also evaluated the winning DL model of the SynthRAD challenge, which used a patch-based approach for centralized MRI-to-sCT [21], namely the MSEP architecture [8]. To enable its use in a FL context, we adapted the original implementation accordingly. To train the MSEP, patches with dimensions $48\times 160\times 160$ were extracted from each sample. The location of each patch was chosen randomly every epoch to increase the amount of possible training input data needed by intrinsic nature of the architecture. Since with this 3D DL approach, the model predicted overlapping patches and the number of predictions was not constant across the volume, the average value was calculated for each voxel using the median voting approach to reach a consensus. Given the computational complexity, the centres included in the federation were reduced to Centre A and Centre C.
The Adam optimizer with a learning rate of $2\times 10^{-4}$ was used. In this scenario, the hardware needed to lead the experiment was an NVIDIA A100 80GB GPU, 32 CPU cores, and 256GB of RAM from a Supermicro AS-4124GS-TNR 4U Rackmount GPU SuperServer.
@@ 4.4.2 Aggregation strategy selection
Further optimisation was pursued through the incorporation of additional techniques, including FedBN and the combination of FedProx and FedBN, with the aim of refining and optimising the federated training process. This was done with a particular focus on ensuring that the global model could converge more rapidly and improve the robustness.
@@ 5.1 Benchmarking deep learning architectures
Table 3 presents the results of the MSEP training conducted in a federated setup. The experiment was completed in a total time of 85 hours and 37 minutes. It illustrates the time required for the federated training and the performance in terms of image similarity. This can be compared with the results obtained from 2D models to determine whether a 3D approach is a suitable candidate for this federated task.
Table 3: Results obtained on the test cases of the local clients (Centre A and C) and on the unseen dataset (Centre E) using the best performing 3D architecture (MSEP) of the SynthRAD2023 challenge for the head MRI to sCT task, as a federated model.
MAE [HU] SSIM PSNR [dB] Total Time
Centre A 229.9 (196.3-232.6) 0.86 (0.86-0.87) 26.56 (26.45-27.83) -
Centre C 219.5 (213.4-231.0) 0.65 (0.59-0.71) 18.98 (18.02-20.09) -
Centre E (Unseen) 240.5 (220.7-250.7) 0.79 (0.77-0.80) 23.70 (23.09-24.29) 85h 37min

### v2

@@ Abstract
The generation of Synthetic Computed Tomography (sCT) images has become a pivotal methodology in modern clinical practice, particularly in the context of Radiotherapy (RT) treatment planning. The use of sCT enables the calculation of doses, pushing towards Magnetic Resonance Imaging (MRI) guided radiotherapy treatments. Moreover, with the introduction of MRI-Positron Emission Tomography (PET) hybrid scanners, the derivation of sCT from MRI can improve the attenuation correction of PET images.
Deep learning methods for MRI-to-sCT have shown promising results, but their reliance on single-centre training dataset limits generalisation capabilities to diverse clinical settings. Moreover, creating centralised multi-centre datasets may pose privacy concerns. To address the aforementioned issues, we introduced FedSynthCT-Brain, an approach based on the Federated Learning (FL) paradigm for MRI-to-sCT in brain imaging. This is among the first applications of FL for MRI-to-sCT, employing a cross-silo horizontal FL approach that allows multiple centres to collaboratively train a U-Net-based deep learning model. We validated our method using real multicentre data from four European and American centres, simulating heterogeneous scanner types and acquisition modalities, and tested its performance on an independent dataset from a centre outside the federation.
In the case of the unseen centre, the federated model achieved a median Mean Absolute Error (MAE) of $102.0$ HU across 23 patients, with an interquartile range of $96.7-110.5$ HU. The median (interquartile range) for the Structural Similarity Index (SSIM) and the Peak Signal to Noise Ratio (PNSR) were $0.89(0.86-0.89)$ and $26.58(25.52-27.42)$, respectively.
The analysis of the results showed acceptable performances of the federated approach, thus highlighting the potential of FL to enhance MRI-to-sCT to improve generalisability and advancing safe and equitable clinical applications while fostering collaboration and preserving data privacy.
@@ 1 Introduction
Intermodality medical Image-To-Image (I2I) translation refers to the process of mapping images from one modality (e.g., Magnetic Resonance Imaging - MRI, Computed Tomography - CT, Positron Emission Tomography - PET) to another, while preserving essential structural and anatomical information. Deep Learning (DL) based techniques have been shown to outperform traditional methods based on image registration and/or voxel-wise analysis [13] and they have gained attention in several clinical fields, the most important being the derivation of Synthetic CT (sCT) from MRI or Cone Beam CT (CBCT), to facilitate radiotherapy planning/adaptation or to improve attenuation correction in PET/MRI acquisition [57, 62, 8].
@@
To improve generalisability in MRI-to-sCT and foster innovation in the field, several studies have been conducted using cross-centre and multi-centre datasets to create models that generalise to prostate and pelvic images from different sources [59, 60, 20]. The study by Han et al. [20] used data from multiple centres also to address the impact of pre-processing to harmonise images and reduce the variability between images using normalisation, clipping and intensity correction techniques. Additionally, it investigated the impact of various combinations of data augmentation techniques for the MRI-to-sCT task focusing on deformation and rotation techniques in the case of multiple centres.
Furthermore, initiatives like the SynthRAD2023 [25] challenge have been proposed by the scientific community. These efforts allow DL models — such as the MSEP framework, the challenge winner [24, 11] — to be trained on multi-centre datasets using both brain and pelvic regions, providing access to more diverse patient data. While directly pooling medical images from multiple sites could intuitively address the issue of small sample sizes, this approach faces significant barriers due to stringent privacy protection policies such as the Health Insurance Portability and Accountability Act (HIPAA)  [^1] https://www.hhs.gov/hipaa/index.html in the United States and the General Data Protection Regulation (GDPR) [^2] https://gdpr-info.eu/ in Europe. These regulations impose limitations on the exchange of personal health data, posing practical challenging to sharing and centralising medical images across sites [29, 19].
@@
To the best of the authors’ knowledge, this is among the first applications in the literature to employ a FL paradigm in MRI-to-sCT. The main contributions and innovative aspects of our work are:
@@
  • The validation of the proposed approach on real multi-centre data, simulating an authentic scenario where each centre handles heterogeneous data due to differences in scanners and acquisition modalities.
  • The testing of the proposed FedSynthCT-Brain on an independent dataset from a centre outside the federation to assess the federated model’s generalization capability.
@@ 2.1 Centralised MRI-to-sCT
Over the past years, numerous DL methods have been implemented for MRI-to-sCT. The methods and models employed differed depending on the target anatomical regions and the computational resources available. Models trained on single image planes (axial, sagittal, or coronal) were termed 2D models. The combination of independently trained 2D models on different planes during or after inference led to improvements in accuracy and was referred to as 2D+ models [57, 8, 13].
Multi-plane models (Multi-2D) employed images from multiple planes to train a single model, leveraging inter-planar information to outperform 2D and 2D+ configurations. Consecutive slices also formed multi-channel input during 2D training, an approach known as 2.5D, which enhanced spatial coherence without requiring full 3D volumes. Smaller image segments were employed across 2D, 2D+, Multi-2D, and 2.5D methods using 2D patching [57, 8, 13].
More recently, the employment of entire 3D volumes or sub-volumes (also named 3D patches) as input provided the most comprehensive spatial information and mitigated inter-slice artefacts observed in 2D-based methods [13]. However, the primary limitation of 3D models lies in their higher computational demand, which requires splitting volumetric data into smaller 3D patches for processing [57]. This patch-based processing may lead to the loss of global context and structural coherence, especially when patches are analysed in isolation [33, 32]. Moreover, the scalability of 3D approaches in FL settings is further constrained due to the increased communication overhead and memory requirements associated with volumetric data. These factors make 3D-DL approaches less suitable for FL real-world scenarios where limited computational resources requirements must be accounted for [3].
The proposed model architectures and methodologies have become increasingly complex in order to achieve accurate results. Among the most widely used models are: U-Nets, Generative Adversarial Networks (GANs), Transformers, and Diffusion-based architectures [57, 8, 13]. In addition to fundamental architectures, recent advancements in the field, such as I2I-Mamba, have demonstrated strong potential for multi-modal medical image synthesis and medical image reconstruction by leveraging state-space operators to enhance contextual interactions across different spatial orientations [5, 28]. Additionally, novel diffusion-based techniques, such as SelfRDB, have been proposed to improve reliability in multi-modal medical image translation, integrating self-consistent recursive sampling and conditional guidance to enhance information transfer between modalities [4].
[…]
Several works employed three different 2-dimensional (2D) fully Convolutional Neural Networks (CNN), to retain structural information in sCTs while simultaneously reducing computational costs, especially when using U-Nets [13, 35, 58, 23]. These 2D methods, by processing slices independently or in parallel, inherently simplify training in distributed environments, such as FL, where data sharing is constrained by computational heterogeneity [3].
@@ 2.2 Federated learning in medical imaging
More recently, FL approaches were proposed for segmentation tasks, such as for brain tumour [63, 2], prostate [74], optic cup and disc segmentation [50, 65]. This led to the development of frameworks that integrate the aforementioned works and are capable of performing segmentations of multiple anatomical structures or different image modalities using FL and foundational models, such as the FedFMS [45], which was created to integrate the Segment Anything Model (SAM) [31] into the medical FL context.
Recent advancements in medical image synthesis have introduced innovative methods, such as FedMed-GAN, which specifically addresses the challenge of synthesizing cross-modality brain MRI images in an FL setting [66]. It employs Federated Averaging (FedAvg), the de facto aggregation strategy algorithm for FL [69], enabling the combination of locally trained models from multiple institutions by computing a weighted average of their parameters.
Inspired by the FL research, this study investigates the feasibility of FL for MRI-to-CT synthesis, addressing a critical gap in the literature. Indeed, to the best of the authors’ knowledge, this is among the first work to apply a cross-silo horizontal FL approach in this context, leveraging multi-center real heterogeneous datasets.
@@ 3.1 Datasets and preprocessing
Parameter Centre A Centre B Centre C Centre D Centre E
Patients # 15 14 21 29 23
Voxel size [mm3] MR 1,1,1 1,1,1 0.78,0.78,1
0.98-1.12,
0.98-1.12,
0.98-1.12
0.98,0.98,0.98
CT
0.49-0.67,
0.49-0.67,
2.5
0.98,
0.98,
3.27
0.78,
0.78,
1
0.69-0.78,
0.69-0.79,
1-3
0.59-1.27,
0.59-1.27,
1-2
Scanners
MR
[T]
MAGNETOM
Trio
[3]
Biograph
mMR
[3]
Vantage
Titan
[1.5]
MAGNETOM
Avanto_fit,
Skyra,
Vida_fit,
Prisma_fit
[1.5-3]
MAGNETOM
Aera,
Avanto_fit
[1.5-3]
CT
[kVp]
LightSpeed
QX/i
[140]
Discovery
ST
[120]
Brilliance
Sensation
Open
[120]
Brilliance
Big
Bore
[120]
SOMATOM
Definition
AS
[120]
Size Axial 256 256 226-357 167-213 167-262
Sagittal 176 176 512 216-262 200-225
Coronal 256 248 512 250-277 225-248
@@
MRIs from Centre B were acquired using a PET-MRI hybrid scanner. Differently from the other centres, this acquisitions were not performed for radiotherapy planning purposes. Thus, the tube energy in the CT was 120kV and the slice thickness was 3.75 mm which introduced partial volume artefacts in the upper part of the skull bone. In light of the substantial difference in the acquisition protocol and scanners, Centre B dataset was used to increase the image variability within the federation (i.e. to introduce noise) to test the robustness of the federated model and simulate a real context.
Images belonging to Centre C presented higher spatial resolution in the cranial caudal direction compared to the other datasets. However, the Centre C’s dataset presented a number of additional challenges, including: the absence of the top of the skull in some patients, the presence of the scanner bed and pillow for each CT, metal artefacts, and high bias field artefact.
The Centre D’s and Centre E’s datasets derived exclusively from two different institutions of the SynthRAD2023 multi-centre dataset (SynthRAD Centre C and SynthRAD Centre B respectively), which was created for the SynthRAD challenge within different MRI-to-sCT algorithms [25]. Before publication, the datasets were cropped to remove the facial region. The FOV covered by both imaging modalities was similar, except for a cut-off at the lower end with a slight angle in addition to a straight horizontal line. In general, Center D and E presented less bias field artefact when compared to Center C. However, between Centre D and E, it was more pronounced in the Centre E dataset. Furthermore, both Centre MRIs were obtained using gadolinium-based contrast agents, thus introducing further heterogeneity within the federation.
To address the image variability, a pre-processing strategy was implemented at each client site with the aim of harmonising data within the federation, without the exchange of data. Following [20], N4 Bias Field Correction [64] was applied to the MRI data to correct intensity inhomogeneity, which can result from factors such as magnetic field variations or patient positioning. Additionally, Min-Max Normalization was applied within each client to reduce variability in pixel intensity values.
@@ 3.3 Deep learning model
To train the aforementioned architecture, each client input 2D slices from each MRI in 3 anatomical planes, axial, coronal, and sagittal, as it is shown in Figure 4. Slices were presented to the network in a randomised order. After extracting all slices from each anatomical plane, they were shuffled to ensure that the DL model received non-sequential input from more than one anatomical plane at a time. This minimised the risk of the model becoming overly dependent on any single plane or orientation, encouraging comprehensive and plane-invariant feature learning across all planes [30]. This method was adapted from the Multi-2D approach [57], modified to enhance generalisation. As a result, it was renamed Random Multi-2D.
@@ 4.3 Performance metrics
In order to evaluate the efficacy of the federated model both on the clients test datasets (centres A, B, C, and D), and unseen test dataset (centre E) standard image similarity metrics in the MRI to CT translation task were employed [57, 8]. These included:
@@ 4.4.1 Deep learning architecture selection
We explored different DL architectures used for MRI-to-sCT conversion, including Conditional GAN, CycleGAN and Transformers (see the supplementary A for further details). Among these, we further investigated several variants of the U-Net architecture which has been used for centralised MRI-to-sCT-Simple U-Net [52], Spadea et al. U-Net [58], Fu et al. U-Net [16], Li et al. U-Nets [37]. These variants included variations in implementations, such as changes in network depth, feature map sizes, and the integration of attention mechanisms. Our analysis focused on evaluating the balance between computational complexity and generalization capability across these architectures.
An additional investigation is presented in B which involved a comparison of the adopted training methodology (Random Multi-2D) with other methodologies (such as 2D+, 2D patches and classic Multi-2D) using the U-Net presented in Section 3.3.
@@ 4.4.2 Aggregation strategy selection
Further optimisation was pursued through the incorporation of additional techniques, including FedBN [40], and the combination of FedProx and FedBN, with the aim of refining and optimising the federated training process. In our study, the FedBN method was specifically designed by the exclusion of batch normalization layers from communication between the server and client. This was done with a particular focus on ensuring that the global model could converge more rapidly and improve the robustness.
@@ 5.1 Benchmarking deep learning architectures
Additional results obtained with different architectures, which proved to be less suitable than U-Nets for the proposed federated task, are presented in Table 4. Nevertheless, the results obtained from the supplementary investigation on the performance obtained of the U-Net by Li et al. [41] when using different DL training methodologies are reported in Table 5.